How much coffee costing $3 a pound should be mixed with 5 pounds of coffee costing $3.50 a pound obtain a mixture costing $3.25 a pound?
Answer by mananth(16946)

coffee $3 a pound x pounds
5 pounds of coffee costing $3.50
mixture costing $3.25 a pound -----x+5 pounds
...
3x+3.50*5=3.25(x+5)
3x+17.50=3.25x+16.25
3.25x-3x=17.50-16.25
0.25x=1.25
/0.25
x=1.25/0.25
x= 5 pounds of $ 3 coffee
...m.annath@hotmail.ca